OpenStudy (chillhill):

Each day, a factory recorded the number (x) of boxes it produces and the total production cost (y) dollars. Equation of the regression: y=10.7x+121. The factory sells the boxes for $19.99 each. Find the least number of boxes that the factory should produce in one day in order to make a profit.

rebeccaxhawaii (rebeccaxhawaii):

10.7 is the price to make?

OpenStudy (chillhill):

I don't know, that's all the information I was given

OpenStudy (kobeni-chan):

I think there's a certain way to do this, but I can't think of how. I do know that you can fill in a guess for the variable and work from there. If x = 10, then the equation would be 10.7(10) + 121 = 107 + 121 = 228 y = 228, and if boxes sell for 19.99 each, then selling 10 boxes would make 199.9 = 228

OpenStudy (kobeni-chan):

These aren't equal, but the numbers aren't too far off from being so. You can try solving the equation with another number, larger or smaller, and see if you can get the numbers closer.